# Interaction induced reentrance of Bose glass and quench dynamics of Bose gases in twisted bilayer and quasicrystal optical lattices

### v1

@@ Appendix C Time dependent Gutzwiller approach
[…]
@@
$-J\left[C_{i,n+1}\sqrt{n_{l}+1}\left(\sum_{j}\Phi_{j}^{*}\right)\right.$ (4)
$\left.+C_{i,n-1}\sqrt{n_{l}}\left(\sum_{j}\Phi_{j}\right)\right]$
@@
Since the time-dependent Gutzwiller method effectively conserves particle number, we can numerically simulate the system’s evolution once the equations of motion for the wavefunction coefficients are obtained. By analyzing typical real-space distributions at different times, we can investigate the system’s dynamical characteristics. This can be illustrated by considering an arbitrary initial Gutzwiller state $\left|\textrm{GW}\left(0\right)\right\rangle$. After a time step $dt$, the state becomes:
$\left|\textrm{GW}\left(dt\right)\right\rangle=\left|\textrm{GW}\left(0\right)\right\rangle-idt\hat{H}\left|\textrm{GW}\left(0\right)\right\rangle$ (5)
The expected value of the particle number at the next time step is:
$\left\langle\hat{N}\left(dt\right)\right\rangle$ $=\left\langle\textrm{GW}\left(dt\right)\right|\hat{N}\left|\textrm{GW}\left(dt\right)\right\rangle$
$=\left\langle\hat{N}\left(0\right)\right\rangle+idt\left\langle\textrm{GW}\left(0\right)\right|\left[\hat{H},\hat{N}\right]\left|\textrm{GW}\left(0\right)\right\rangle$ (6)
@@
The fact that particle number is not exactly conserved, but only on average, is reflected in the commutator of the second term, which does not vanish exactly, but whose expectation value is zero for any arbitrary state.

### v3

@@ Appendix C Time dependent Gutzwiller approach
In this appendix, we present the technical details of the time-dependent Gutzwiller approach [78, 79, 80]. This approach is built on top of the static Gutzwiller approach. The static Gutzwiller method is employed to determine the ground state by minimizing the total energy of the system, given by $E\left(\left\{C_{i,n}\right\}\right)=\left\langle\textrm{GW}\right|\hat{H}\left|\textrm{GW}\right\rangle$, while the dynamic Gutzwiller method minimizes the action to simulate the system’s evolution. Specifically, by applying the variational principle, we can write the action of the system as $S=\int dt\left[-i\left\langle\textrm{GW}\left(t\right)\right|\partial_{t}\left|\textrm{GW}\left(t\right)\right\rangle+\left\langle\textrm{GW}\left(t\right)\right|\hat{H}\left|\textrm{GW}\left(t\right)\right\rangle\right]$. Variation with respect to the coefficients $C_{i,n}\left(t\right)$ yields the equation of motion:
@@
$-J\left[C_{i,n+1}\sqrt{n+1}\left(\sum_{j}\Phi_{j}^{*}\right)\right$. (5)
$\left.+C_{i,n-1}\sqrt{n}\left(\sum_{j}\Phi_{j}\right)\right]$
@@
Time-dependent Gutzwiller method automatically conserves the expectation value total particle number $\left\langle\hat{N}\right\rangle$ as long as $\hat{H},\hat{N}]=0$. This can be illustrated by considering a generic Gutzwiller state at time $t\left|\textrm{GW}\left(t\right)\right\rangle$. After a time step $dt$, the state becomes:
$\left|\textrm{GW}\left(t+dt\right)\right\rangle=\left|\textrm{GW}\left(t\right)\right\rangle-idt\hat{H}\left|\textrm{GW}\left(t\right)\right\rangle$ (6)
The expectation value of the total particle number at the next time step is
$\left\langle\hat{N}\left(t+dt\right)\right\rangle$ $=\left\langle\textrm{GW}\left(t+dt\right)\right|\hat{N}\left|\textrm{GW}\left(t+dt\right)\right\rangle$
$=\left\langle\hat{N}\left(t\right)\right\rangle+idt\left\langle\textrm{GW}\left(t\right)\right|\left[\hat{H},\hat{N}\right]\left|\textrm{GW}\left(t\right)\right\rangle$ (7)
@@
indicating $\left\langle\hat{N}\right\rangle$ is conserved during evolution as long as $\hat{H},\hat{N}]=0$.